Preliminary list of participants International Colloquium on Combinatorics and Graph Theory July 15-19, Balatonlelle, Hungary

D. Acketa, Institute of Mathematics, TRG D. Obradovica 4, 21000 Novi Sad, Yugoslavia W 9:00 acketa@unsim.ns.ac.yu G. Alexe, Universitatea din Galati, Catedra de matematica, str. Domneasca nr. 47, 6200 Galati, Romania galex@lgal.sfos.ro F 11:20 N. Alon, Dept. of Mathematics, Tel Aviv University, Ramat Aviv 69978, Israel noga@math.tau.ac.il Tu 9:00 J.L. Arocha, IMATE-UNAM, Ciudad Universitaria, Area de la Ivestig. Cientifica, Circuito Exterior, Mexico D.F., Mexico arocha@concha.matem.unam.mx Tu 18:30 B. Bajnok, Dept. of Math. and Comp. Sci., Gettysburg College, Gettysburg, PA 17325, United States of America bbajnok@gettysburg.edu G. Bacsó, ELTE TTK Számítógéptud. Tsz., H-1088, Budapest, Múzeum krt. 6-8., K.T. Balinska, Technical University of Poznan, Computer Science Center, Pl. M. Sklodowskiej-Curie 5, 60-965 Poznan, Poland balinska@ruby.kari.put.poznan.pl W 12:25 J. Balogh, JATE Bolyai Intézet, H-6720, Szeged, Aradi vértanúk tere 1. jobal@sol.cc.u-szeged.hu F 10:55 J. Baltzer, Baltzer Science Publishers, Asterweg 1A, 1031 HL Amsterdam laurenz@ns.baltzer.nl H.-J. Bandelt, Math. Seminar, Universität Hamburg, Bundesstr. 55, D-20146 Hamburg, Germany J. Baráth, JATE Bolyai Intézet, H-6720, Szeged, Aradi vértanúk tere 1. O. Barr, Math. Department, Umea University, 901 87 Umea, Sweden barr@abeel.math.umu.se Th 11:20 I. Bárány, MTA MKI, H-1053, Budapest, Reáltanoda u. 13-15., barany@math-inst.hu I. Beck, College of Stavanger, Dept. of Mathematics, P. Box 2557 Ullandhaug, 4004 Stavanger, Norway istvan.beck@tn.his.no S. Bezrukov, Dept. of Mathematics and Comp. Sci., University of Paderborn, Warburger Str. 100, D-33102 Paderborn, Germany sb@uni-paderborn.de Tu 11:20 B. Bollobás, Memphis State University, Memphis, Tennessee, US; University of Cambridge, Cambridge, UK bollobab@ibex.msci.memphis.edu bollobas@pmms.cam.ac.uk Th 15:30 P. Brass, Universität Greifswald, Institut für Mathematik, Jahnstrasse 15a, D-17487 Greifswald, Germany brass@math-inf.uni-greifswald.d400.de Tu 10:55 Y. Caro, Dept. of Mathematics, School of Education, Univ. of Haifa-Oranim, Tivon 36910, Israel zeac603@uvm.haifa.ac.il M. Chari, Department of Mathematics, Louisiana State University, Baton Rouge, LA 70803, United States of America chari@math.lsu.edu Th 15:55 G. Csizmadia, H-1119, Budapest, Fehérvári út 183. csizi@cs.elte.hu Tu 15:30 W. Deuber, Fakultät für Mathematik, Universität Bielefeld, Postfach 100131, D-33501 Bielefeld, Germany deuber@mathematik.uni-bielefeld.de M 15:30 A. Dress, Univ. Bielefeld, Postfach 10 01 31, D-33501 Bielefeld, Germany Y. Egawa, Department of Applied Mathematics, Science University of Tokyo, 1-3 Kagurazaka, Shinjuku-ku, Tokyo 162, Japan Th 12:10 G. Elekes, ELTE TTK Számítógéptud. Tsz. H-1088, Budapest, Múzeum krt. 6-8. elekes@cs.elte.hu M 15:55 S. Eliahou, University of Geneva, Dept. of Mathematics, 2-4 rue de Lievre, CP 240, 1211 Geneva 24, Switzerland shalom@sc2a.unige.ch M 16:20 K. Engel, Universität Rostock, FB Mathematik, PF 999, D-18051 Rostock, Germany konrad.engel@mathematik.uni-rostock.d400 Th 11:45 P. Erdős, MTA MKI, H-1053, Budapest, Reáltanoda u. 13-15. M 10:30 P.L. Erdős, MTA MKI, H-1053, Budapest, Reáltanoda u. 13-15. elp@math-inst.hu Tu 10:30, Th 15:55 R. Faudree, Dept. of Math. Sciences, Memphis State Univ., Memphis, TN 38152, United States of America faudreer@memstuxi.bitnet Th 10:30 F. Fodor, H-5600, Békéscsaba, Lencsési út 7.I.3. M. Forys, Jagellonian University, Inst. of Comp. Sci., Nawojki 11, 30-072 Krakow, Poland forysm@softlab.ii.uj.edu.pl F 18:30 W. Forys, Jagellonian University, Inst. of Comp. Sci., Nawojki 11, 30-072 Krakow, Poland forysw@softlab.ii.uj.edu.pl A. Frolov, Mathematical Modeling Department, Moscow Power Eng. Institute, Krasnokazarmennaya 14, 111250 Moscow, Russia frolov@apmsun.mpei.ac.ru M 18:30 Z. Füredi, MTA MKI, H-1053, Budapest, Reáltanoda u. 13-15. furedi@math-inst.hu A. Gál, School of Mathematics, Institute for Advanced Study, Olden Lane, Princeton, NJ 08540, US panni@math.ias.edu W 10:45 H. Galeana-Sanchez, Instituto de Matematicas, UNAM, Circuito Exterior, Ciudad Universitaria, Mexico 04510, D.F., Mexico edicion@gauss.matem.unam.mx Th 18:05 A. Geroldinger, Inst. f. Mathematik, Univ. Graz, Heinrichstrasse 36, A-8010 Graz, Austria geroldin@bedvgm.kfunigraz.ac.at M 16:45 E.Kh. Gimadi, Institute of Mathematics of RAS, Universitetskii prosp. 4, 630090 Novosibirsk 90, Russia gimadi@math.nsk.su J. Goldman, JATE Bolyai Intézet, H-6720, Szeged, Aradi vértanúk tere 1. h531818@stud.u-szeged.hu J.R. Griggs, Department of Mathematics, University of South Carolina, Columbia, SC 29208, US griggs@math.sc.edu Tu 17:40 H. Gropp, Mühlingstr. 19, D-69121 Heidelberg, Germany d12@ix.urz.uni-heidelberg.de W 10:45 G. Gutin, Dept. of Math. and Comp. Sci., Odense University, 5230 Odense M, Denmark gutin@imada.ou.dk W 11:10 A. Gyárfás, MTA SzTAKI, H-1111, Budapest, Lágymányosi u. 11 gyarfas@luna.aszi.sztaki.hu E. Győri, MTA MKI, H-1053, Budapest, Reáltanoda u. 13-15. ervin@math-inst.hu R. Häggkvist Dept. of Math., Umea University, S-901 87 Umea, Sweden rolandh@vimse.umdc.umu.se Th 16:45 P. Hajnal, JATE Bolyai Intézet, H-6720, Szeged, Aradi vértanúk tere 1. hajnal@math.u-szeged.hu J. Harant, Technische Hochschule Ilmenau, Inst. für Mathematik, D-98693 Ilmenau, Germany harant@mathematik.tu-ilmenau.de Tu 17:40 S. Hartmann, Universität Rostock, FB Mathematik, D-18051 Rostock, Germany sven@zeus.math.uni-rostock.de W 9:25 I. Havel, Mathematical Institute, Czech Academy of Sciences, ul. Zitna 25, 11567 Praha 1, Czech Republic haveli@beba.cesnet.cz Th 17:40 P.E. Haxell, Dept. of Combinatorics and Optimization, University of Waterloo, Waterloo, ONT N2L 3G1, Canada pehaxell@math.uwaterloo.ca Tu 11:20 T. Hegedűs, Dept. of Comp. Sci., Comenius University, 84215 Bratislava, Slovakia hegedus@fmph.uniba.sk P. Hlineny, Charles University, Dept. of Appl. Mathematics, Malostranske nam. 25, 118 00 Praha 1, Czech Republic hlineny@kam.ms.mff.cuni.cz Tu 16:45 M. Hogh, Abteilung für Theor. Biologie, Kirschallee 1, D-53115 Bonn, Germany hogh@sn-bota-1.bota.uni-bonn.de V.P. Il’ev, P.O. Box 8501, Omsk 644070, Russia iljev@univer.omsk.su C. Iliopoulos, King’s College London, Department of Computer Science, London WC2R 2LS, U. K. csi@dcs.kcl.ac.uk Th 16:20 É. Jakó, ELTE TTK, H-1088, Budapest, Múzeum krt. 6-8 jako@ludens.elte.hu J. Jegier, Inst. of Mathematics, University of Mining and M., Al. Mickiewicza 30, 30059 Krakow, Poland jegiw@uci.agh.edu.pl S. Jendrol, Dept.of Geom.and Algebra, P.J.Safarik University, Jesenna 5, 041 54 Kosice, Slovakia jendrol@kosice.upjs.sk M 15:30 T. Jordán, BME Mat. Szám.tud. Tsz., H-1111, Budapest, Egry J. u. 20-22. jordan@cs.elte.hu M. Juvan, University of Ljubljana, Department of Mathematics, Jadranska 19, 61111 Ljubljana, Slovenia martin.juvan@fmf.uni-lj.si Tu 16:20

D. Kaller, School of Computing Science, Simon Fraser University, Burnaby, B.C. V5A 1S6, Canada kaller@cs.sfu.ca

Gy. Katona, MTA MKI, H-1053, Budapest, Reáltanoda u. 13-15. ohkatona@math-inst.hu

Y. Gy. Katona, MTA MKI, H-1053, Budapest, Reáltanoda u. 13-15. kiskat@cs.elte.hu Th 16:20

J.C.M. Keijsper, Fakulteit WINS, Univ. of Amsterdam, Plantage Muidergracht 24, 1018 TV Amsterdam, The Netherlands keijsper@fwi.uva.nl

L.H. Khachatrian, Fakultät für Mathematik, Universität Bielefeld, Postfach 100131, D-33501 Bielefeld Germany lk@mathematik.uni-bielefeld.de Tu 10:30

Z. Király, ELTE TTK Számítógéptud. Tsz., H-1088, Budapest, Múzeum krt. 6-8., kiraly@cs.elte.hu F 12:10

Á. Kisvölcsey, MTA MKI, H-1053, Budapest, Reáltanoda u. 13-15 ksvlcs@math-inst.hu

S. Klavzar, Dept. of Mathematics, PEF, University of Maribor, Koroska cesta 160, 2000 Maribor, Slovenia sandi.klavzar@uni-lj.si Tu 12:10

M. Klazar, KAM, Malostranske nam. 25, 118 00 Praha 1, Czech Republic klazar@kam.ms.mff.cuni.cz Th 16:45

V.F. Kolchin, Steklov Mathematical Institute, 42, ul. Vavilova, Moscow 117966, Russia kolchin@mian.su Tu 15:30

A.A. Kolokolov, Inst. of Info.Techn. and Appl. Math., 28, ul. Andrianova, Omsk 644077, Russia kolo@iitam.omsk.su F 17:40

P. Komjáth, ELTE TTK Számítógéptud. Tsz., H-1088, Budapest, Múzeum krt. 6-8. kope@cs.elte.hu M 18:30

A.V. Kostochka, Intstitute of Mathematics, Universitetskiy pr. 4, 630090 Novosibirsk, Russia sasha@math.nsk.su M 16:20

J. Kratochvil, KAM MFF UK (Charles Univ.), Malostranske nam 25, 118 00 Praha 1, Czech Republic honza@kam.ms.mff.cuni.cz Th 10:30 A. Kuribayashi, Dept. of Mathematics, Chuo University, 1-13-27, Kasuga Bunkyo-ku, Tokyo 112, Japan kuri@math.chuo-u.ac.jp Th 12:10

N.N. Kuzjurin, Institute for System Programming, Russian Academy of Sciences, B. Communisticheskaja 25, 109004 Moscow, Russia nnkuz@ivann.delta.msk.su Th 16:20

R. Labahn, Universität Rostock, FB Mathematik, D-18051 Rostock, Germany roger.labahn@mathematik.uni-rostock.de F 17:40

U. Leck, Universität Rostock, FB Mathematik, D-18051 Rostock, Germany uleck@zeus.math.uni-rostock.de Th 10:30

L. Lipták, H-6722, Szeged, Mérey u. 20/B.I.8., liptak@csilla.inf.u-szeged.hu

L. Lovász, ELTE TTK Számítógéptud. Tsz., H-1088, Budapest, Múzeum krt. 6-8., lovasz@cs.elte.hu

D.D. Lozovanu, Institute of Mathematics, Academy of Sciences, str. Academiei 5, 277028 Kishinev, Moldova 17ddl@mathem. moldova.su F 18:05

T. Luczak, Institute of Math., Adam Mickiewicz Univ., Matejki 48/49, 60-769 Poznan, Poland Tu 10:30

A. Lukács, ELTE TTK Számítógéptud. Tsz., H-1088, Budapest, Múzeum krt. 6-8. lukacs@cs.elte.hu

P.H. Lundow, Umea University, Department of Mathematics, 901 87 Umea, Sweden phl@abel.math.umu.se

A.A. Makhnev, Institute of Math. and Mech., UB RAS, Kovalevskoi 16, 620219 Ekaterinburg, Russia mak@top.imm.intec.ru Th 11:20

U. Matthias, MTA MKI, H-1053 Budapest, Reáltanoda u. 13-15 matthias@math-inst.hu

F. Matus, Universität Bielefeld, Fakultät für Wirtschaftswissenschaften, Statistik und Informatik, Postfach 100131, D-33501 Bielefeld, Germany mfrantisek@wiwi.uni-bielefeld.de Th 15:30

C. Mauduit, Institut de Mathem. Discretes, UPR 9016, case 930, 163 Av. de Luminy, 13288 Marseille Cedex 9, France mauduit@iml.univ-mrs.fr M 11:40 P.G. Mezey, Univ. of Saskatchewan, Mathematical Chemistry, 110 Science Place, Saskatoon, SK S7N 5C9, Canada mezey@sask.usask.ca Tu 11:20

P. Mihók, KGA PF UPJS, Jesenna 5, 04154 Kosice, Slovakia mihok@kosice.upjs.sk M 18:30

B. Mohar, Dept. of Maths., University of Ljubljana, Jadranska 19, 61111 Ljubljana, Slovenia bojan.mohar@uni-lj.si Tu 15:55

M. Molloy, Department of Computer Science, University of Toronto, Toronto, ONT M5S 3G4, Canada molloy@cs.toronto.edu M 18:05

L. Morales, Univ. of Texas at Dallas, Computer Science Program, EC 31, P.O. Box 830688, Richardson, TX 75083-0688, United States of America morales@utdallas.edu

K. Murota, RIMS, Kyoto University, Japan murota@kurims.kyoto-u.ac.jp F 10:30

H. Müller, Friedrich-Schiller Universität, Institut für Infomatik, D-07740 Jena, Germany hm@minet.uni-jena.de Th 10:55

E. Nanchen, University of Geneva, Dept. of Mathematics, 2-4 rue de Lievre, C.P. 240, 1211 Geneva 24, Switzerland nanchen@divsun.unige.ch Th 18:30

J. Nesetril, KAM, MFF UK, Charles University, Malostranske nam 25, 11800 Praha 1, Czech Republic, nesetril@cspoguk11.bitnet W 9:00

E. Olaru, Universitatea din Galati, Catedra de matematica, str. Domneasca nr. 47, 6220 Galati, Romania eolaru@ugal.ro F 10:30

J. Opatrny, Dept. of Comp. Sci., Concordia University, 1455 de Maisonneuve w., Montreal H3G 1M8, Canada opatrny@cs.concordia.ca F 15:30

J.G. Oxley, Dept. of Mathematics, Louisiana State University, Baton Rouge,LA 70803-4918, US oxley@math.lsu.edu F 11:20

V.I. Pagurova, Dept. of Comp. Math. and Cyb., Moscow State University, Vorobiovy Gory, 119899 Moscow, Russia pvi@cmc.msu.su Tu 16:20

R.A. Pendavingh, Fakulteit WINS, Univ. of Amsterdam, Plantage Muidergracht 24, 1018 TV Amsterdam, The Netherlands pendavin@fwi.uva.nl F 18:05 E. Perekhodtseva, Vernadsky av. 95-1-112, Moscow, Russia dima@dmp.fpfe.mpt.ru M 16:45 G. Pete, H-6726, Szeged, Középfasor 15/c. h432427@stud.u-szeged.hu P. Pitteloud, University of Geneva, Dept. of Mathematics, 2-4 rue de Lievre, CP 240, 1211 Geneva 24, Switzerland pittelou@ibm.unige.ch H.-J. Prömel, Sektion Mathematik, Humboldt Universität, Berlin, Germany proemel@informatik.hu-berlin.de F 15:30 L.V. Quintas, Mathematics Dept., Pace University, New York, NY 10038, US quintas@pacevm.dac.pace.edu W 12:00 S.B. Rao, Stat. Math. Division, Indian Statistical Inst., 203 Barrackpore Trunk Road, Calcutta 700035, India. Th 15:30 A. Recski, BME Villamosmérn.Kar Mat.Tsz. , H-1111, Budapest, Műegyetem rakpart 1. recski@vma.bme.hu F.S. Roberts DIMACS, Rutgers University, P.O. Box 1179, Piscataway NJ 08855-1179, US froberts@dimacs.rutgers.edu Tu 15:30 Y. Roditty, School of Math. Sci., Dept. of Comp. Sci., Tel Aviv Univ., Ramat-Aviv, Tel Aviv 69978, Israel jr@math.tau.ac.il M. Rosenfeld, Dept. of Math. GN-50, Univ. of Washington, Seattle, WA 98195, US moishe@math.washington.edu Th 10:55 V. Rödl, Dept. Math., Emory University, Atlanta, GA 30322, US rodl@mathcs.emory.edu Th 9:00 M. Ruszinkó, MTA SzTAKI, H-1111, Budapest, Kende u. 13-17. ruszinko@lutra.sztaki.hu Tu 12:10 A. Rycerz, Institute of Mathematics, Univ. of Mining and Metallurgy, Al. Mickiewicza 30, 30-059 Krakow, Poland grrycerz@cyf-kr.edu.pl E. Sampathkumar, Department of Mathematics, University of Mysore, Mysore-570006, India protein@cscftri.nic.in F 18:30 A.A. Sapozhenko, Moscow State University, Dept. of Comp. Math. and. Cybernetics, Vorobiowy Gory, 119899 Moscow, Russia mathcyb@cs.msu.su Th 11:20 G. Sárközy, 209 South 33rd Street, Philadelphia, PA 19104-6395, US gsarkozy@math.upeenn.edu Tu 11:45 N. Sauer, Dept. of Math. and Statistics, University of Calgary, 2500 Univ. Drive N. W., Calgary, Alberta T2N 1N4, Canada nsauer@acs.ucalgary.ca M 17:40 R. Schneider, Geblerg. 18/3, A-1170 Wien, Austria Rudolf.R.Schneider@siemens.at J. Schönheim, School of Math. Sci., Tel-Aviv University, Tel Aviv, Ramat-Aviv 69978, Israel meira@math.tau.ac.il A. Schrijver, CWI, P.O. Box 4079, 1009 AB Amsterdam, The Netherlands lex@cwi.nl F 9:00 I. Sciriha, c/o ’Troika’, G. Montebebello Str., Tarxien PLA 12, Malta irene@jaguar.is.unimt.mt F 16:20 A.P. Sergeyev, Borodyanka, Parkowaya 4/48, 255770 Kiev Region, Ukraine post@sveng.kiev.ua W 11:10 S. Simic, Department of Mathematics, Faculty of Electrical Engineering, University of Belgrade, Bulevar Revolucije 73, 11000 Belgrade, Yugoslavia esimics@ubbg.etf.bg.ac.yu F 16:45 G. Simonyi MTA MKI, H-1053, Budapest, Reáltanoda u. 13-15. simonyi@math-inst.hu Tu 11:45 M. Simonovits MTA MKI, H-1053, Budapest, Reáltanoda u. 13-15. miki@math-inst.hu V.G. Skobelev, Inst. of Appl. Math. and Mech., NAS of Ukraine, R. Luxemburg Str., 74, 340114 Donetsk, Ukarine leonid@dsam.donetsk.ua W 12:00 R. Skrekovski, Department of Mathematics, University of Ljubljana, Jadranska 19, 61111 Ljubljana, Slovenia skreko@fmf.uni-lj.si M 17:40 T. Slivnik, DPMMS, 16 Mill Lane, Cambridge, U. K. T.Slivnik@pmms.cam.ac.uk W.F. Smyth, Dept. of Comp. Sci. and Syst., McMaster University, 1280 Main Str. W., Hamilton, Ontario L8S 4K1, Canada smyth@mcmaster.ca M 17:40, 18:05 D. Solomon, Institute of Mathematics, Academy of Sciences, str. Academiei 5, 277028 Kishinev, Moldova 17ddl@mathem.moldova.su F 18:05 B.S. Stechkin, Steklov Mathematical Institute, 42, ul. Vavilova, Moscow 117966, Russia stechkin@log.mian.su G. Steiner, McMaster University, MGD-415, 1280 Main Street West, Hamilton, ONT L8S 4M4, Canada steiner@mcmaster.ca W 11:35 B. Sudakov, School of Mathematical Sciences, Tel Aviv University, Ramat Aviv, Tel Aviv 69978, Israel sudakov@math.tau.ac.il M 15:55 H. Sudborough, Computer Science Program, University of Texas at Dallas, Richardson, Texas 75083-0688, US hal@utdallas.edu M 16:20 I. Szalkai, Veszprémi Egyetem Mat. Tsz., H-8200, Veszprém, Egyetem tér 10. szalkai@almos.vein.hu F 15:55 L. Székely, ELTE TTK Analízis Tanszék, H-1088, Budapest, Múzeum krt. 6-8., szekely@konig.elte.hu Tu 10:55 T. Szkaliczki, BME Mat. és Szám.tud. Tanszék H-1521, Budapest, Egry J. u. 20-22. H ép. III. sztibor@vma.bme.hu F 16:45 T. Szőnyi, ELTE TTK Számítógéptud. Tsz., H-1088, Budapest, Múzeum krt. 6-8., szonyi@cs.elte.hu Th 11:45 B. Toft, Mathematics and Computer Science, Odense University, 5230 Odense M, Denmark btoft@imada.ou.dk M 11:40 S. Todorcevic, Th 17:40 Zs. Tuza, MTA SzTAKI, H-1111, Budapest, Kende u. 13-17. tuza@lutra.sztaki.hu R. Urbaniak, Konrad-Zuse-Zentrum für Informationstechnik Berlin, Heilbronner Str. 10, D-10711 Berlin, Germany urbaniak@zib-berlin.de F 16:20 P. Valencia Saravia, Tu 18:30 B. Velickovic, Dept. of Mathematics, Univ. of California, Berkeley, CA 94720, United States of America. Th 18:05 A. Vdovina, Kamenetska 80, apt. 15, 280000 Khmelnitsky, Ukraine bob@ceko.freenet.kiev.ua Tu 18:05

A. Vesel, Dept. of Mathematics, PEF, University of Maribor, Koroska cesta 160, 2000 Maribor, Slovenia vesel@uni-mb.si Tu 12:10

K. Vesztergombi, H-1125, Budapest, Kútvölgyi út 71.

A. von Haeseler, Zool. Institute, Univ. of Munich, Luisenstr. 14, D-80021 Munich, Germany arndt@zi.biologie.uni-muenchen.de M 11:40

A. Wagler, Konrad-Zuse-Zentrum für Informationstechnik Berlin, Heilbronner Str. 10, D-10711 Berlin, Germany wagler@ZIB-berlin.de F 16:20

T. Warnow, Univ. of Pennsylvania, Dept. of Comp. Sci., Philadelphia 19104, United States of America tandy@central.cis.upenn.edu M 15:30

N. Zagaglia Salvi, Dip. di Matematica, Politecnico di Milano, P.za L. da Vinci 32, 20133 Milano, Italy norzag@mate.polimi.it W 11:35

A. Zaks, 3 Keereen Kayemet Str., Herzelia 46406, Israel ayalz@math.tau.ac.il F 12:10

A.M. Zubkov, Steklov Mathematical Institute, 42, ul. Vavilova, Moscow 117966, Russia Tu 15:55

Yu.A. Zuev, Computing Center Russian Academy of Sciences, Vavilova 40, 117967 Moscow, Russia zuev@ccas.ru Tu 16:45